OpenStudy (anonymous):

-9x^2-12x+4=49 solve by using square root property

OpenStudy (lgbasallote):

turn it into a square of a binomial \[(-3x + 2)^2 = 49\] get the square root of both sides \[-3x + 2 = \pm 7\]

OpenStudy (lgbasallote):

\[-3x = \pm 7 - 2\] \[-3x = 5; -3x = -9\] can you take it from there?

OpenStudy (anonymous):

 I think it is +9x^2